OpenStudy (anonymous):

Need help working out f(x) problem Given: f(x) = 2x + 5 and g(x) = x² and h(x) = -2x find h(g(f(x))) = Am I supposed to find f(x) first?

OpenStudy (anonymous):

Take f(x) and plug into g(x) first g(x)=x^2 g(f(x)) = (2x+5)^2 Now plug that into h(x) h(x) = -2x h(g(f(x))) = -2(2x+5)^2